OpenStudy (anonymous):

PLEASE HELP…AND EXPLAIN write the point-slope form of an equation for a line that passes through the point with the given slope.(3,3), m=-1

OpenStudy (anonymous):

do you know what point-slope form looks like?

OpenStudy (anonymous):

would it be a or b a: y-3=-1-(x+3) B: y-3=-1(x-3)

OpenStudy (anonymous):

y-y1=m(x-x1) all you have to do is plug it in

OpenStudy (anonymous):

and yes its, y-y1=m(x-x1)

OpenStudy (anonymous):

it's b, plug 3,3 into x1, y1 and plug -1 into m (the slope) so yes b